# Random Sperner lemma and random Brouwer fixed point theorem

### v3

@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
    $G$ is said to be finitely stable if $I_{a}x+I_{a^{c}}y\in G$ for any $x,y\in G$ and any $a\in B_{\mathcal{F}}$. $G$ is said to be $\sigma$-stable (or to have the countable concatenation property in terms of [23]) if, for each sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and each $\{a_{n},n\in\mathbb{N}\}\in p(1)$, there exists some $x\in G$ such that $I_{a_{n}}x=I_{a_{n}}x_{n}$ for each $n\in\mathbb{N}$, such an $x$ is unique since $E$ is assumed to be regular, usually denoted by $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$, called the countable concatenation of $\{x_{n},n\in\mathbb{N}\}$ along $\{a_{n},n\in\mathbb{N}\}$. By the way, if $G$ is $\sigma$-stable and $H$ is a nonempty subset of $G$, then $\sigma(H):=\{\sum_{n=1}^{\infty}I_{a_{n}}h_{n}:\{h_{n},n\in\mathbb{N}\}$ is a sequence in $H$ and $\{a_{n},n\in\mathbb{N}\}\in p(1)\}$ is called the $\sigma$-stable hull of $H$.
For any $(n-1)$-dimensional $L^{0}$-simplex $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ in an $L^{0}$-module $E$, as shown by Guo et al. in [26], $S$ is $\sigma$-stable and each element $x\in S$ can be uniquely represented as $x=\sum^{n}_{i=1}\xi_{i}x_{i}$, where $\xi_{i}\in L^{0}_{+}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ and are such that $\sum^{n}_{i=1}\xi_{i}=1$. Moreover, since $E$ is also a linear space over $\mathbb{K}$ and the $L^{0}$-affine independence of $\{x_{1},x_{2},\cdots,x_{n}\}$ implies their affine independence, then $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ is a classical $(n-1)$-dimensional simplex in the linear space $E$, where $Conv(G)$ denotes the usual convex hull of a nonempty subset $G$ in a linear space. In this case, we call $S^{\prime}$ the counterpart simplex of $S$. Throughout this paper, whenever we refer to an $L^{0}$-simplex $S$, we will always denote its counterpart simplex by $S^{\prime}$.
Recall from [1] that for any $(n-1)$-dimensional simplex $,S^{\prime}=Conv(\{x_{1},x_{2},$$\cdots,x_{n}\})$ in a linear space, a finite family $\mathscr{S}^{\prime}$ of $(n-1)$-dimensional simplexes is called a simplicial subdivision of $S^{\prime}$ if
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$. A finite family $\mathscr{S}$ of $(n-1)$-dimensional $L^{0}$-simplexes is called an $L^{0}$-simplicial subdivision of $S$ if
@@
Let $\mathscr{S}$ and $\mathscr{S}^{\prime}$ be the same as in Definition 2.6, then we call $\mathscr{S}^{\prime}$ the counterpart simplicial subdivision of $\mathscr{S}$. Throughout this paper, whenever we refer to an $L^{0}$-simplicial subdivision $\mathscr{S}$, we will always denote its counterpart simplicial subdivision by $\mathscr{S}^{\prime}$. $L^{0}$-simplexes are much more complicated objects than usual simplexes, we encounter lots of challenges in the initial study of them, these force us to introduce Definition 2.6 in connection with the classical simplicial subdivision so that we can make full use of the theory of classical simplicial subdivisions to give a thorough treatment of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes.
The notion of $L^{0}$-extreme points was introduced and studied by Guo et al. in [26]. For two nonempty subsets $G$ and $H$ of an $L^{0}$-module (resp., a linear sapce) such that $H\subseteq G$, $H$ is called an $L^{0}$-extreme subset (resp., an extreme subset) of $G$ if both $x$ and $y$ belong to $H$ whenever $x$ and $y$ belong to $G$ and there exists $\lambda\in L^{0}(\mathcal{F},(0,1))$ (resp., $\lambda\in(0,1)$) such that $\lambda x+(1-\lambda)y\in H$. Further, if $H=\{h\}$ is a singleton and an $L^{0}$-extreme subset (resp., an extreme subset) of $G$, then $h$ is called an $L^{0}$-extreme point (resp., an extreme point) of $G$. We will always denote by $ext_{L^{0}}(G)$ (resp., $ext(G)$) the set of $L^{0}$-extreme points (resp., extreme points) of $G$. For any family $\mathscr{G}$ of nonempty subsets of an $L^{0}$-module (resp., a linear space), the set of $L^{0}$-extreme points (resp., extreme points) of $\mathscr{G}$ is defined by $ext_{L^{0}}(\mathscr{G}):=\sigma[\bigcup_{G\in\mathscr{G}}ext_{L^{0}}(G)]$ (resp., $ext(\mathscr{G}):=\bigcup_{G\in\mathscr{G}}ext(G)$).
For any $(n-1)$-dimensional simplex $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ in a linear space and any simplicial subdivision $\mathscr{S}^{\prime}$ of $S^{\prime}$, a mapping $\psi:ext(\mathscr{S}^{\prime})\rightarrow\{1,2,\cdots,n\}$ is called a proper labeling function of $\mathscr{S}^{\prime}$ (see [1]) if
@@
where $\chi(y):=\{i:\alpha_{i}>0\}$ for any $y=\sum_{i=1}^{n}\alpha_{i}x_{i}$. In fact, the set of proper labeling functions is exactly $\Pi_{y\in ext(\mathscr{S}^{\prime})}\chi(y)$.
@@
In [11, Definition 2.1], Drapeau et al. introduced the notion of a proper $L^{0}$-labeling function for the $L^{0}$-barycentric subdivision of an $L^{0}$-simplex in $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Definition 2.7 below is a generalization of [11, Definition 2.1]. But we adopt a strengthened notion of a completely labeled $L^{0}$-simplex since such a notion is both easy to understand and such a completely labeled $L^{0}$-simplex indeed exists (see Theorem 2.9). In fact, according to Definition 2.7, $\phi(y_{j})$ is an equivalence class of a constant valued random variable for each $j\in\{1,2,\cdots,n\}$ for a completely labeled $L^{0}$-simplex $G=Conv_{L^{0}}(\{y_{1},y_{2},\cdots,y_{n}\})$, and hence our notion of a completely labeled $L^{0}$-simplex is completely similar to the notion of a completely labeled simplex.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$, $\mathscr{S}$ an $L^{0}$-simplicial subdivision of $S$ and $\phi:ext_{L^{0}}(\mathscr{S})\rightarrow L^{0}(\mathcal{F},\{1,2,\cdots,n\})$ a mapping. $\phi$ is called an $L^{0}$-labeling function of $\mathscr{S}$ if $\phi$ is $\sigma$-stable. An $L^{0}$-labeling function $\phi$ of $\mathscr{S}$ is said to be proper if
@@
The classical Sperner lemma, originally established by Sperner in [48] was formulated for simplexes in $\mathbb{R}^{n}$. Let $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional simplex in a linear space $L$ and $L^{n}=Span(\{x_{1},x_{2},\cdots,x_{n}\})$ be the linear subspace of $L$ generated by $\{x_{1},x_{2},\cdots,x_{n}\}$, then $L^{n}$ is at most $n$-dimensional and the classical Sperner lemma is only related to $L^{n}$ and is independent of $L$, so one can readily see the following version of Sperner lemma for simplexes in a general linear space.
@@
Let $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional simplex in a linear space, $\mathscr{S}^{\prime}$ a simplicial subdivision of $S^{\prime}$ and $\psi$ a proper labeling function of $\mathscr{S}^{\prime}$. Then there are an odd number of completely labeled simplexes in $\mathscr{S}^{\prime}$
For any given sequence $\{G_{n},n\in\mathbb{N}\}$ of nonempty subsets of a $\sigma$-stable set $G$ and any given $\{a_{n},n\in\mathbb{N}\}\in p(1)$,
@@
Theorem 2.9 below can be aptly called the random Sperner lemma.
Theorem 2.9.
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$, $\mathscr{S}$ an $L^{0}$-simplicial subdivision of $S$ and $\phi$ a proper $L^{0}$-labeling function of $\mathscr{S}$. Then there exist a finite family $\{S_{m}:=Conv_{L^{0}}(\{y_{1}^{m},y_{2}^{m},\cdots,y_{n}^{m}\}),m=1\sim M\}$ of $(n-1)$-dimensional $L^{0}$-simplexes in $\mathscr{S}$ and $\{a_{m},m=1\sim M\}\in p(1)$ such that
@@
is an $(n-1)$-dimensional $L^{0}$-simplex completely labeled by $\phi$ and satisfies that
@@
In Theorem 2.9, $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ is an $(n-1)$-dimensional $L^{0}$-simplex contained in $S$ such that each vertex of $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ belongs to $ext_{L^{0}}(\mathscr{S})$ since both $S$ and $ext_{L^{0}}(\mathscr{S})$ are $\sigma$-stable. Besides, $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ is not necessarily in $\mathscr{S}$ but merely a finite concatenation of the finite subfamily $\{S_{m}:m=1\sim M\}$ of $\mathscr{S}$, which is a significant difference between Theorem 2.9 and the classical Sperner lemma (namely Lemma 2.8).
3 Proof of Theorem 2.9
We first give the following two Examples — Examples 3.1 and 3.2. The first one shows the $L^{0}$-barycentric subdivision introduced in [11, Definition 1.10] is a special case of our Definition 2.6. The second one shows our Definition 2.6 does not just include the $L^{0}$-barycentric subdivision.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$ and $G(n)$ be the group of permutations of $\{1,2,\cdots,n\}$. For any $\pi\in G(n)$, define
@@
Then, by applying the same method as in [11, Lemma 1.11], one can verify that $\{S_{\pi}:\pi\in G(n)\}$ is an $L^{0}$-simplicial subdivision of $S$, which is called the $L^{0}$-barycentric subdivision of $S$.
@@
(1). On the one hand, for any $x=\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}x_{3}$ in $S$, where $\lambda_{1},\lambda_{2},\lambda_{3}\in L_{+}^{0}(\mathcal{F})$ and $\lambda_{1}+\lambda_{2}+\lambda_{3}=1$, let
@@
which implies that $S\subseteq\sigma\{\bigcup\mathscr{S}\}$. On the other hand, since $S$ is both $\sigma$-stable and $L^{0}$-convex, it follows that $\sigma\{\bigcup\mathscr{S}\}\subseteq S$. Therefore, $S=\sigma\{\bigcup\mathscr{S}\}$.
(2). For any $z\in Conv_{L^{0}}(\{x_{1},y_{1,2},y_{1,3}\})\cap Conv_{L^{0}}(\{y_{1,2},y_{1,3},y_{2,3}\})$, let
@@
[…]
@@
[…]
@@
(3). Similar to the argument presented in (2), one can readily verify that
@@
To prove Theorem 2.9, we need the following Lemma 3.3, Lemma 3.4 and Theorem 3.5. Although (2) of Lemma 3.3 is almost clear since each $S_{i}$ is $\sigma$-stable, we still give its proof for the reader’s convenience.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$. Then we have the following statements:
@@
    For any sequence $\{S^{m}=Conv_{L^{0}}(\{x_{1}^{m},x_{2}^{m},\cdots,x_{n}^{m}\}),m\in\mathbb{N}\}$ of $(n-1)$-dimensional $L^{0}$-simplexes contained in $S$ and any $\{a_{m},m\in\mathbb{N}\}\in p(1)$, $\sum_{m=1}^{\infty}I_{a_{m}}S^{m}$ is an $(n-1)$-dimensional $L^{0}$-simplex contained in $S$ with vertexes $\sum_{m=1}^{\infty}I_{a_{m}}x_{1}^{m},\sum_{m=1}^{\infty}I_{a_{m}}x_{2}^{m},\cdots,$ $\sum_{m=1}^{\infty}I_{a_{m}}x_{n}^{m}$.
@@
(1). First, we prove that
@@
and $S^{m}$ is an $(n-1)$-dimensional $L^{0}$-simplex, then
@@
(2). Since $S$ is $\sigma$-stable, for any $\{b_{i},i=1\sim I\}\in p(1)$, we have
@@
$b_{i}=\bigvee_{n\in N_{i}}a_{n},\forall i\in\{1,2,\cdots,I\},$
@@
As a direct application of Lemma 5.4 of [26], Lemma 3.4 below gives an interesting connection between the set of $L^{0}$-extreme points of an $L^{0}$-simplicial subdivision and the set of extreme points of its counterpart simplicial subdivision.
@@
Let $E$ and $S$ be the same as in Lemma 3.3, and $\mathscr{S}=\{S_{i},i=1\sim I\}$ be an $L^{0}$-simplicial subdivision of $S$. Then we have the following statements:
@@
(2). Since $ext_{L^{0}}(S_{i})=\sigma(ext(S^{\prime}_{i}))$ for each $i\in\{1,2,\cdots,I\}$, we have
@@
Throughout this paper, for any $\xi=\sum_{i=1}^{n}I_{a_{i}}i\in L^{0}(\mathcal{F},\{1,2,\cdots,n\})$, in such an expression of $\xi$, from now on we always assume that $a_{i}=(\xi=i)$ for each $i\in\{1,2,\cdots,n\}$ (then it is clear that $\{a_{i},i=1\sim n\}\in p(1)$), we will also always use $\mathcal{N}_{\xi}$ for the set $\{i:a_{i}>0\}$.

### v5

@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
    $G$ is said to be finitely stable if $I_{a}x+I_{a^{c}}y\in G$ for any $x,y\in G$ and any $a\in B_{\mathcal{F}}$. $G$ is said to be $\sigma$-stable (or to have the countable concatenation property in terms of [24]) if, for each sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and each $\{a_{n},n\in\mathbb{N}\}\in p(1)$, there exists some $x\in G$ such that $I_{a_{n}}x=I_{a_{n}}x_{n}$ for each $n\in\mathbb{N}$. Such an $x$ is unique since $E$ is assumed to be regular, usually denoted by $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$, called the countable concatenation of $\{x_{n},n\in\mathbb{N}\}$ along $\{a_{n},n\in\mathbb{N}\}$. If $G$ is $\sigma$-stable and $H$ is a nonempty subset of $G$, then $\sigma(H):=\{\sum_{n=1}^{\infty}I_{a_{n}}h_{n}:\{h_{n},n\in\mathbb{N}\}$ is a sequence in $H$ and $\{a_{n},n\in\mathbb{N}\}\in p(1)\}$ is called the $\sigma$-stable hull of $H$.
For any (n1)(n-1)-dimensional $L^{0}$-simplex $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ in an $L^{0}$-module $E$, as shown by Guo et al. [27], $S$ is $\sigma$-stable and each element $x\in S$ can be uniquely represented as $x=\sum^{n}_{i=1}\xi_{i}x_{i}$, where $\xi_{i}\in L^{0}_{+}(\mathcal{F})$ for any $i\in\{1,2,\cdots,n\}$ such that $\sum^{n}_{i=1}\xi_{i}=1$. Moreover, since $E$ is also a linear space over $\mathbb{K}$ and the $L^{0}$-affine independence of $\{x_{1},x_{2},\cdots,x_{n}\}$ implies their affine independence, then $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ is a classical (n1)(n-1)-dimensional simplex in the linear space $E$, where $Conv(G)$ denotes the usual convex hull of a nonempty subset $G$ in a linear space. In this case, we call $S^{\prime}$ the associated simplex of $S$. Throughout this paper, whenever we refer to an $L^{0}$-simplex $S$, we will always denote its associated simplex by $S^{\prime}$.
Recall from [1] that for any (n1)(n-1)-dimensional simplex $S^{\prime}=Conv(\{x_{1},x_{2},$ $\cdots,x_{n}\})$ in a linear space, a finite family $\mathscr{S}^{\prime}$ of (n1)(n-1)-dimensional simplexes is called a simplicial subdivision of $S^{\prime}$ if
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an (n1)(n-1)-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$. A finite family $\mathscr{S}$ of (n1)(n-1)-dimensional $L^{0}$-simplexes is called an $L^{0}$-simplicial subdivision of $S$ if
@@
Let $\mathscr{S}$ and $\mathscr{S}^{\prime}$ be the same as in Definition 2.6, then we call $\mathscr{S}^{\prime}$ the associated simplicial subdivision of $\mathscr{S}$. Throughout this paper, whenever we refer to an $L^{0}$-simplicial subdivision $\mathscr{S}$, we will always denote its associated simplicial subdivision by $\mathscr{S}^{\prime}$. $L^{0}$-simplexes are much more complicated objects than usual simplexes, causing considerable challenges in our initial study of them. These challenges motivate us to introduce Definition 2.6 in connection with the classical simplicial subdivision so that we can make full use of the theory of classical simplicial subdivisions to give a thorough treatment of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes.
The notion of $L^{0}$-extreme points was first introduced and studied by Guo et al. in [27]. For two nonempty subsets $G$ and $H$ of an $L^{0}$-module (resp., a linear space) such that $\subseteq G. HH$ is called an $L^{0}$-extreme subset (resp., an extreme subset) of $G$ if both $x$ and $y$ belong to $H$ whenever $x$ and $y$ belong to $G$ and there exists $\lambda\in L^{0}(\mathcal{F},(0,1))$ (resp., $\lambda\in(0,1)$) such that $\lambda x+(1-\lambda)y\in H$. Further, if $H=\{h\}$ is a singleton and an $L^{0}$-extreme subset (resp., an extreme subset) of $G$, then $h$ is called an $L^{0}$-extreme point (resp., an extreme point) of $G$. We will always denote by $ext_{L^{0}}(G)$ (resp., $ext(G)$) the set of $L^{0}$-extreme points (resp., extreme points) of $G$. For any family $\mathscr{G}$ of nonempty subsets of an $L^{0}$-module (resp., a linear space), the set of $L^{0}$-extreme points (resp., extreme points) of $\mathscr{G}$ is defined by $ext_{L^{0}}(\mathscr{G}):=\sigma[\bigcup_{G\in\mathscr{G}}ext_{L^{0}}(G)]$ (resp., $ext(\mathscr{G}):=\bigcup_{G\in\mathscr{G}}ext(G)$).
For any (n1)(n-1)-dimensional simplex $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ in a linear space and any simplicial subdivision $\mathscr{S}^{\prime}$ of $S^{\prime}$, a mapping $\psi:ext(\mathscr{S}^{\prime})\rightarrow\{1,2,\cdots,n\}$ is called a proper labeling function of $\mathscr{S}^{\prime}$ (see [1]) if
@@
where $\chi(y):=\{i:\alpha_{i}>0\}$ for any $y=\sum_{i=1}^{n}\alpha_{i}x_{i}$. In fact, the set of all proper labeling functions is exactly $\Pi_{y\in ext(\mathscr{S}^{\prime})}\chi(y)$.
@@
In [12, Definition 2.1], Drapeau et al. introduced the notion of a proper $L^{0}$-labeling function for the $L^{0}$-barycentric subdivision of an $L^{0}$-simplex in $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Definition 2.7 below is a generalization of [12, Definition 2.1]. But we adopt a strengthened notion of a completely labeled $L^{0}$-simplex since such a notion is both easy to understand and such a completely labeled $L^{0}$-simplex indeed exists (see Theorem 2.9). In fact, according to Definition 2.7, $\phi(y_{j})$ is an equivalence class of a constant valued random variable for each $j\in\{1,2,\cdots,n\}$ for a completely labeled $L^{0}$-simplex $G=Conv_{L^{0}}(\{y_{1},y_{2},\cdots,y_{n}\})$, and hence our notion of a completely labeled $L^{0}$-simplex is completely similar to the notion of a completely labeled simplex.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an (n1)(n-1)-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$, $\mathscr{S}$ an $L^{0}$-simplicial subdivision of $S$ and $\phi:ext_{L^{0}}(\mathscr{S})\rightarrow L^{0}(\mathcal{F},\{1,2,\cdots,n\})$. The mapping $\phi$ is called an $L^{0}$-labeling function of $\mathscr{S}$ if $\phi$ is $\sigma$-stable. An $L^{0}$-labeling function $\phi$ of $\mathscr{S}$ is said to be proper if
@@
The classical Sperner lemma, originally established by Sperner in [51] was formulated for simplexes in $\mathbb{R}^{n}$. Let $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ be an (n1)(n-1)-dimensional simplex in a linear space $L$ and $L^{n}=Span(\{x_{1},x_{2},\cdots,x_{n}\})$ be the linear subspace of $L$ generated by $\{x_{1},x_{2},\cdots,x_{n}\}$. Then $L^{n}$ is at most $n$-dimensional and the classical Sperner lemma is only related to $L^{n}$ and is independent of $L$, so one can readily see the following version of Sperner lemma for simplexes in a general linear space.
@@
Let $S^{\prime}=Conv(\{x_{1},x_{2},\cdots,x_{n}\})$ be an (n1)(n-1)-dimensional simplex in a linear space, $\mathscr{S}^{\prime}$ a simplicial subdivision of $S^{\prime}$ and $\psi$ a proper labeling function of $\mathscr{S}^{\prime}$. Then there are an odd number of completely labeled simplexes in $\mathscr{S}^{\prime}$ .
For any given sequence $\{G_{n},n\in\mathbb{N}\}$ of nonempty subsets of a $\sigma$-stable set $G$ of an $L^{0}$-module and any given $\{a_{n},n\in\mathbb{N}\}\in p(1)$,
@@
Theorem 2.9 below can be aptly called a random Sperner lemma.
Theorem 2.9 (Random Sperner Lemma).
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an (n1)(n-1)-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$, $\mathscr{S}$ an $L^{0}$-simplicial subdivision of $S$ and $\phi$ a proper $L^{0}$-labeling function of $\mathscr{S}$. Then there exist a finite family $\{S_{m}:=Conv_{L^{0}}(\{y_{1}^{m},y_{2}^{m},\cdots,y_{n}^{m}\}),m=1\sim M\}$ of (n1)(n-1)-dimensional $L^{0}$-simplexes in $\mathscr{S}$ and $\{a_{m},m=1\sim M\}\in p(1)$ such that
@@
is an (n1)(n-1)-dimensional $L^{0}$-simplex completely labeled by $\phi$ and satisfies that
@@
In Theorem 2.9, $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ is an (n1)(n-1)-dimensional $L^{0}$-simplex contained in $S$ such that each vertex of $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ belongs to $ext_{L^{0}}(\mathscr{S})$ since both $S$ and $ext_{L^{0}}(\mathscr{S})$ are $\sigma$-stable. Besides, $\sum_{m=1}^{M}I_{a_{m}}S_{m}$ is not necessarily in $\mathscr{S}$ but merely a finite concatenation of the finite subfamily $\{S_{m}:m=1\sim M\}$ of $\mathscr{S}$, which is a significant difference between Theorem 2.9 and the classical Sperner lemma (namely, Lemma 2.8).
3 Proof of Theorem 2.9: Random Sperner Lemma
We first give the following two examples — Examples 3.1 and 3.2. The first one shows the $L^{0}$-barycentric subdivision introduced in [12, Definition 1.10] is a special case of our Definition 2.6. The second one shows that our Definition 2.6 surpasses the $L^{0}$-barycentric subdivision.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an (n1)(n-1)-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$ and $G(n)$ be the group of permutations of $\{1,2,\cdots,n\}$. For any $\pi\in G(n)$, define
@@
Then, by applying the same method as in [12, Lemma 1.11], one can verify that $\{S_{\pi}:\pi\in G(n)\}$ is an $L^{0}$-simplicial subdivision of $S$, which is called the $L^{0}$-barycentric subdivision of $S$.
@@
(1) On the one hand, for any $x=\lambda_{1}x_{1}+\lambda_{2}x_{2}+\lambda_{3}x_{3}$ in $S$, where $\lambda_{1},\lambda_{2},\lambda_{3}\in L_{+}^{0}(\mathcal{F})$ and $\lambda_{1}+\lambda_{2}+\lambda_{3}=1$, let
@@
implying $S\subseteq\sigma\{\bigcup\mathscr{S}\}$. On the other hand, since $S$ is both $\sigma$-stable and $L^{0}$-convex, it follows that $\sigma\{\bigcup\mathscr{S}\}\subseteq S$. Therefore, $S=\sigma\{\bigcup\mathscr{S}\}$.
(2) For any $z\in Conv_{L^{0}}(\{x_{1},y_{1,2},y_{1,3}\})\cap Conv_{L^{0}}(\{y_{1,2},y_{1,3},y_{2,3}\})$, let
@@
$\begin{cases}\begin{aligned} \xi_{1}+\frac{\xi_{2}}{2}+\frac{\xi_{3}}{2}&=\frac{\eta_{2}+\eta_{3}}{2}\\ \frac{\xi_{2}}{2}&=\frac{\eta_{1}+\eta_{3}}{2}\\ \frac{\xi_{3}}{2}&=\frac{\eta_{1}+\eta_{2}}{2}\end{aligned}\end{cases},$
@@
$\begin{cases}\begin{aligned} \xi_{1}&=\eta_{1}=0\\ \xi_{2}&=\eta_{3}\\ \xi_{3}&=\eta_{2}\end{aligned}\end{cases},$
@@
(3) Similar to the argument presented in (2), one can readily verify that
@@
To prove Theorem 2.9, we need the following Lemma 3.3, Lemma 3.4 and Theorem 3.5. Although part (2) of Lemma 3.3 is almost clear since each $S_{i}$ is $\sigma$-stable, we still give its proof here for the reader’s convenience.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an (n1)(n-1)-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$. Then we have the following properties:
@@
    For any sequence $\{S^{m}=Conv_{L^{0}}(\{x_{1}^{m},x_{2}^{m},\cdots,x_{n}^{m}\}),m\in\mathbb{N}\}$ of (n1)(n-1)-dimensional $L^{0}$-simplexes contained in $S$ and any $\{a_{m},m\in\mathbb{N}\}\in p(1)$, $\sum_{m=1}^{\infty}I_{a_{m}}S^{m}$ is an (n1)(n-1)-dimensional $L^{0}$-simplex contained in $S$ with vertexes $\sum_{m=1}^{\infty}I_{a_{m}}x_{1}^{m},\sum_{m=1}^{\infty}I_{a_{m}}x_{2}^{m},\cdots,$ $\sum_{m=1}^{\infty}I_{a_{m}}x_{n}^{m}$.
@@
(1) First, we prove that
@@
and $S^{m}$ is an (n1)(n-1)-dimensional $L^{0}$-simplex, we have
@@
(2) Since $S$ is $\sigma$-stable, for any $\{b_{i},i=1\sim I\}\in p(1)$, we have
@@
$b_{i}=\bigvee_{n\in N_{i}}a_{n},\forall i\in\{1,2,\cdots,I\}.$
@@
As a direct application of Lemma 5.4 of [27], Lemma 3.4 below gives an interesting connection between the set of $L^{0}$-extreme points of an $L^{0}$-simplicial subdivision and the set of extreme points of its associated simplicial subdivision.
@@
Let $E$ and $S$ be the same as in Lemma 3.3, and $\mathscr{S}=\{S_{i},i=1\sim I\}$ be an $L^{0}$-simplicial subdivision of $S$. Then we have the following relations:
@@
(2) Since $ext_{L^{0}}(S_{i})=\sigma(ext(S^{\prime}_{i}))$ for each $i\in\{1,2,\cdots,I\}$, we have
@@
For any $\xi=\sum_{i=1}^{n}I_{a_{i}}i\in L^{0}(\mathcal{F},\{1,2,\cdots,n\})$, in such an expression of $\xi$, from now on we always assume that $a_{i}=(\xi=i)$ for each $i\in\{1,2,\cdots,n\}$ (then it is clear that $\{a_{i},i=1\sim n\}\in p(1)$). We will also always use $\mathcal{N}_{\xi}$ for the set $\{i:a_{i}>0\}$.